OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = x + 3 and the given point (5,0).

OpenStudy (anonymous):

first, what's the slope of the given line, y = x + 3 ???

OpenStudy (anonymous):

is it 1

OpenStudy (anonymous):

yes... so m = 1 and we have a point it goes through, (5, 0). do you know point-slope form ?

OpenStudy (anonymous):

or how about just slope-intercept form, y=mx + b ???

OpenStudy (anonymous):

point-slope form: \(\large y-y_1=m(x-x_1) \), where m is the slope, \(\large (x_1,y_1) \) is the known point slope-intercept form: \(\large y=mx+b \), where m is the slope, b is the y-intercept

OpenStudy (anonymous):

so the equation would be y=5/0x+1

OpenStudy (anonymous):

The equation would be \[y-0=1(x-5)\]

OpenStudy (anonymous):

the slope of the given eq. is 1.. by slope intercept form we have y−y1=m(x−x1), now substitute m=1, y1=0, x1=5 the required eq. is y-0=1(x-5) or x-y=5